OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = 11 –4x + 2y = 2 Answer and formula or just answer need fast :)

OpenStudy (anonymous):

#bohotness

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

ansers (1, 3) (2, 3) (1, 2) (1, –2)

OpenStudy (bohotness):

\(\huge~good~job\) \(\huge~welcome~to~openstudys\)

OpenStudy (anonymous):

thanks:P

OpenStudy (bohotness):

your welcoem

OpenStudy (bohotness):

which one you think it is

OpenStudy (anonymous):

i have no clue but i was only half way through the lesson but i got to finish it soon becuase if i dont well i wont have high enough % to get a a- or b not sure which it is

OpenStudy (bohotness):

okay

OpenStudy (bohotness):

2x + 3y = 11 2( 2x + 3y = 11) 4x + 6y = 22 –4x + 2y = 2 -4x + 2y = 2 8y = 24 y = 3

OpenStudy (anonymous):

ok now i just put it in to th equation

OpenStudy (bohotness):

yes :_)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

nvm thats wrong its not on teh answers

OpenStudy (anonymous):

sinmplifie maby so 2,3

OpenStudy (bohotness):

try b

OpenStudy (anonymous):

k

OpenStudy (anonymous):

another one?

OpenStudy (bohotness):

going head

OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 8g – 5h = 7 4g + 10h = 1

OpenStudy (anonymous):

?=-

OpenStudy (bohotness):

try post it again>

OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 8g - 5h = 7 4g + 10h = 1 (Points : 1)

OpenStudy (bohotness):

?

OpenStudy (anonymous):

k whatever ill work it out myself

OpenStudy (bohotness):

whats your answer choices

OpenStudy (anonymous):

nvm, i already did thanks for earlyer help getting off now its late so peace :)

OpenStudy (bohotness):

bye:)

OpenStudy (anonymous):

I realydon't know

OpenStudy (bohotness):

okay lets start kk